Please answer all questions at the end of the expl
Answer #1

53. answer: C

removal of nH3 shifts equilibrium to right , so that NH3 production increases.

54. by increasing ksp solubility increases.

so that ,

answer: b

55. [H3O+] = 10^-14 / (10^-4.223) = 1.67*10^-10 M

answer: c

56. answer: a

a weakacid and its salt with satrongbase

57. pH of weak monoproticacid = 1/2(pka-logC)

     1.98 = 1/2(x-log0.6)

   x = pka = 3.74

   ka = 10^-3.74

       = 1.82*10^-4

answer: C

58 . answer: A